Ticket Number: HEA-NOC/19981216-01 Ticket Status: UPDATE
Ticket Type: Unscheduled Ticket Source: HEA-NOC
Ticket Scope: Line Site/Line: HEAnet <==> Janet
Ticket Owner: [log in to unmask] Problem Fixer: TE/BT/JANET
Ticket Opened: Problem Start: 19981216 14:25 UTC
Ticket Closed: Problem End:
The is a problem with the relability of the line between HEAnet and
Queen's University, Belfast.
All traffic to and from JANET, TEN-34 and TEN-155.
Traffic is being routed via Cable & Wireless in the U.S.
Problem has been reported to T.E. Ticket #352.
19981216 17:45. T.E. have tested and found no problems. Search has
been transferred to Janet NOSC and Belfast. Both HEA-NOC and Janet
NOSC have tried unsuccessfully to get intervention at Belfast. Line
is presently up but still unacceptably error-prone.
19981216 18:20. BT have been contacted and will arrange for testing
of line from 09:00am to 10:00am tomorrow morning at Queen's University,
Belfast. Line is presently up but still showing errors.
19981217 13:40. NOC have looped the line & B.T. are testing in Belfast.
Currently running clean; test will continue for 15 mins.
19981218 16:00. Circuit currently clean; neither BT nor TE took action.
Will monitor until tomorrow morning & reevaluate with NOSC.
Time to Fix:
Clean since 15:00